y = sqrt(x+1)^2 - sqrt(x-1)^2

Answer by checkley71(8403)

Y=SQRT(X+1)^2-SQRT(X-1)^2
Y=X+1-(X-1)
Y=X+1-X+1
Y=2 ANSWER.
2=SQRT(X+1)^2-SQRT(X-1)^2
2=(X+1)-(X-1)
2=X+1-X+1
2=2 THUS X=0